# Uncovering Latent Arguments in Social Media Messaging by Employing LLMs-in-the-Loop Strategy

### v1

@@ A.3.1 Climate Campaigns
CustBasedAltEnergy ”Customer engagement in alternative energy solutions, emphasizing renewable and nuclear options to combat climate change and promote sustainability.”, ”Public support and awareness of nuclear energy as a sustainable, carbon-free alternative energy solution in the U.S.”, ”Community engagement and support for wind energy, highlighting its benefits for jobs, rural economies, and sustainable development.”, ”Community engagement in renewable energy through a cost-effective, equipment-free Community Solar program with enrollment incentives.”, ”Homeowner participation in solar energy trials to foster renewable energy adoption and reduce individual energy costs.”, ”Embracing solar energy for its broad benefits, focusing on economic growth, environmental protection, and enhanced accessibility.”, ”The adoption and support of utility-scale solar projects to supply clean energy to the grid for consumers.”, ”Embracing wind and nuclear power as key to Illinois’ energy strategy, highlighting job creation and environmental benefits.”, ”Adopting biofuels and innovative fuel technologies to reduce emissions and promote sustainability in transportation for consumers.”, ”Community engagement and legal actions to promote sustainable development and energy practices at the local level.”, ”Adopting geothermal heating and cooling as a cost-effective, efficient, and environmentally friendly alternative energy solution for homeowners.”
@@
FoodSecurity ”Addressing climate change’s impact on agriculture, social equity, and women’s rights to ensure food security for all.”, ”Cross-sector collaboration, addressing food loss, engaging communities, and prioritizing food security initiatives to combat global hunger.”, ”Participation in the Ration Challenge to support emergency food aid and attending a forum to explore food security issues.”, ”Engagement with the Fruitful Communities initiative to combat climate change, promote sustainable agriculture, and enhance food security.”, ”Urgent humanitarian aid and community support to address immediate and long-term food security challenges in crisis-affected areas.”, ”Sustainable dietary changes, particularly reducing meat consumption, to improve environmental health and ensure long-term food security.”, ”Community support and mobilization to provide essential resources and aid, ensuring food security and health during crises.”, ”Combining sustainable agriculture, traditional knowledge, and community efforts to improve food security and environmental sustainability.”
@@
EnrgyAffrd&SustainLeg ”Legislative action in Virginia to ensure energy reliability and affordability, learning from Texas’ deregulation disaster.”, ”Legislative action to significantly lower energy costs, promoting cleaner and more reliable energy solutions in North Carolina.”, ”Legislative and policy efforts to enhance energy efficiency, grid reliability, and sustainability across various regions.”, ”Legislative supporting utility bill assistance, pollution reduction, energy-efficient building codes, and transition to pollution-free buildings for sustainability.”, ”Policy-driven support for renewable energy, affordability, emissions reduction, and sustainable development, emphasizing community involvement and legislative backing.”, ”Legislative action to ensure utility service protection, improve energy infrastructure resilience, and address affordability amid extreme weather and economic challenges.”, ”Transitioning to sustainable energy sources, enhancing security for clean energy businesses, and supporting policies for cleaner, affordable energy solutions.”, ”Leveraging legislation and community programs to provide financial assistance for utility bills, ensuring energy affordability for all residents.”
@@
EcofrndlyConsmChoice ”Making eco-friendly consumer choices by purchasing sustainable apparel and actively contributing to environmental conservation efforts.”, ”Choosing eco-friendly, sustainable laundry solutions like Earthbreeze to reduce environmental impact and support ethical consumption.”, ”Adopting home water treatment systems to reduce reliance on single-use plastic bottles, promoting environmental sustainability and health.”, ”Making eco-friendly consumer choices by using Cleanyst’s plant-based products to reduce plastic waste and carbon emissions.”, ”Empowering communities and individuals through education and involvement in sustainable practices to combat climate change effectively.”, ”Embracing sustainable living through eco-friendly consumer choices, corporate sustainability efforts, and supporting green energy and conservation initiatives.”, ”Choosing sustainable, skin-friendly products like bamboo bandages as a simple yet effective way to support eco-friendly consumer choices.”, ”Supporting local leadership and community initiatives that promote sustainability, equitable growth, and eco-friendly consumer practices in urban development.”
@@
PlasticWste&EnvImpact ”Recognizing plastic’s sustainability benefits, urging scientific studies and informed policies to optimize its use and minimize environmental impact.”, ”Reducing plastic waste through reusable products, supporting cleanup efforts, and embracing innovative solutions to protect marine ecosystems.”, ”Using pyrolysis technology to recycle waste plastic, supporting a circular economy and environmental protection efforts.”, ”Urgent collective action, habit changes, and support for organizations fighting plastic pollution to protect oceans and marine life.”, ”Eengaging brands towards Plastic Neutrality and supports innovative recycling technologies to combat the global plastic waste crisis.”, ”Daily action against plastic waste, supporting legislation, and making conscious choices in food consumption and waste management.”, ”Responsible waste management, recycling efforts, and addressing illegal activities to mitigate environmental contamination and promote sustainability.”
@@
PromtSustainTrnsport ”Embracing sustainable transportation and clean energy through EV test drives and educational events on innovative energy solutions.”, ”Attending Drive Electric Earth Day events to learn about and embrace the benefits of electric vehicles for a sustainable future.”, ”Adopting cycling and supporting policies for electric school buses and a bold climate bill for sustainable transportation.”, ”Urgent investment in electric vehicles and legislative support to reduce emissions and create union jobs.”, ”Adopting electric vehicles for their environmental benefits, cost savings, and the expanding availability of models.”, ”Adopting clean car standards and investing in EV infrastructure to improve air quality, health, and combat climate change.”, ”Significant investment in public transportation electrification and infrastructure to support climate goals, community benefits, and union job creation.”, ”Adopting and promoting sustainable transportation solutions and technologies to improve air quality and address environmental challenges.”, ”Infrastructure investments and community initiatives that reduce emissions and promote sustainable transportation as essential to combating climate change.”, ”Adopting biking, walking, electric vehicles, and public transit as key strategies to reduce emissions and improve air quality.”, ”Shifting to public and active transportation, electric vehicles, and renewable energy to reduce emissions and enhance sustainability.”
@@
WaterMngmnt&Sustain ”Stricter regulations on coal ash disposal to protect water quality and public health in Georgia.”, ”Supporting legislative efforts to protect and restore Colorado’s rivers for economic growth, climate resilience, and agricultural preservation.”, ”Exploring and adopting successful global strategies to address water scarcity and ensure sustainable water management in rapidly growing states.”, ”Sustainable water management through refurbishing existing facilities and adopting eco-friendly water solutions to protect health and the environment.”, ”Proactive water management strategies, infrastructure investment, and community involvement to ensure sustainable water access and conservation efforts.”, ”Responsible disposal of hazardous wastes like oils and grease to protect water sources and promote environmental sustainability.”, ”Innovative financing, policy reforms, and sustainable practices to ensure water security and resilience against climate change impacts.”, ”Active engagement and informed leadership in water board governance to ensure sustainable water management and protection.”
A.3.2 COVID-19 Vaccine Campaigns
@@
GovDistrust ”Accountability and reform in the Biden administration, criticizing its policies for undermining national stability and public trust.”, ”Against government control of the prescription drug market, warning of hindered medical innovation and reduced access to cures.”, ”Skepticism towards government actions and legislation, emphasizing the need for transparency and resistance to perceived overreach.”, ”Skepticism towards the Biden administration’s handling of COVID-19, immigration, and economic policies.”, ”Questions the efficacy and motives of government policies on immigration, economy, and public health measures.”
@@
GovTrust ”Recognizing and appreciating government efforts and collaboration in pandemic response, emphasizing trust in legislative support for public health.”, ”Recognizing the Biden administration’s efforts to address key national and global issues, building trust through proactive policies and actions.”, ”Political engagement and trust through transparency, community involvement, and proactive governance initiatives.”, ”Building trust through community engagement, open dialogue, and collaboration between citizens and their representatives.”, ”President Biden’s commitment to international accords, public health, economic recovery, and progressive policy achievements to build government trust.”
@@
VaccineRollout ”The widespread and accessible COVID-19 vaccination and testing services, emphasizing no-appointment-necessary clinics to increase public participation.”, ”Utilizing digital platforms to access up-to-date information on COVID-19 vaccine availability and registration across different locations.”, ”Easy and accessible COVID-19 vaccination for all eligible individuals, emphasizing convenience and inclusivity in public health efforts.”, ”The organized distribution of COVID-19 vaccines across Texas, highlighting state efforts to enhance accessibility and efficiency in vaccination.”, ”The importance of local news in keeping communities informed about the evolving COVID-19 vaccine rollout and registration processes.”, ”Accessible vaccine clinics, incentive programs, and safety protocols to enhance vaccine rollout at the local level.”
@@
VaccineSymptom ”Emphasizes the safety and effectiveness of COVID-19 vaccines despite mild side effects, advocating for vaccination to protect against the virus.”, ”Vigilance and thorough investigation into the rare but serious side effects like blood clots associated with the J&J vaccine.”, ”Careful monitoring and transparent communication about potential vaccine side effects to ensure public safety and confidence.”, ”Informed vaccine choices, awareness of rare side effects, and the importance of continued vaccination for public health safety.”, ”Awareness and transparency regarding the range and prevalence of COVID-19 vaccine-related symptoms and effects.”, ”Advocates for booster shots to enhance immunity, addressing vaccine safety and mild side effects experienced post-vaccination.”, ”Advocates for the safety of COVID-19 vaccines for children, emphasizing mild side effects and community protection through vaccination.”
@@
VaccineEquity ”Against price control legislation, emphasizing its potential to hinder vaccine development and equitable access to life-saving COVID-19 vaccinations.”, ”Global vaccine equity, emphasizing free and universal access to COVID-19 vaccines to end the pandemic for all.”, ”Overcoming barriers to ensure equitable COVID-19 vaccine access for underserved communities and prioritizing vulnerable populations.”, ”Investigating and improving the distribution and administration process to ensure equitable vaccine access in Illinois.”, ”Equitable access to COVID-19 vaccines, addressing eligibility, distribution fairness.”, ”Donating funds that will be matched to multiply impact, promoting fair global access to vaccines.”, ”Efforts and challenges in equitable vaccine distribution and access for seniors across various counties.”
@@
VaccineStatus ”Raising awareness about the low vaccination rates in Central Texas counties, as indicated by state data.”, ”Staying informed about COVID-19 vaccination updates and initiatives through trusted local news apps in respective counties.”, ”Accessing COVID-19 vaccine information and locations for vaccination through the SmartNews app for free.”, ”Highlights the concern over the high number of unvaccinated individuals in Central Texas counties according to state data.”
@@
EncourageVaccination ”Promoting vaccination by providing information, guidance, and incentives to encourage people to get vaccinated against COVID-19.”, ”Importance of COVID-19 vaccination by emphasizing its safety, efficacy, availability, and role in ending the pandemic.”, ”Engaging in community discussions and raising awareness about vaccination, especially within the LatinX community.”, ”Encourages COVID-19 vaccination and boosters for all ages to protect against the virus and maintain community health.”, ”Promoting COVID-19 vaccination for children and the community, emphasizing safety, accessibility, and the benefits of vaccination.”
@@
VaccineMandate ”Against COVID vaccine mandates and passports, urging support for personal choice and constitutional rights regarding vaccination.”, ”Awareness and updates on issues related to COVID-19 vaccine mandates and refusals in various states.”, ”COVID-19 vaccine mandates, including recent Supreme Court rulings and public opinion on federal mandates.”, ”Against Biden’s vaccine mandate, portraying it as government tyranny and urging resistance and financial support.”, ”Advocating for Covid-19 vaccine mandates for customers by Snowy Range Taxi in Wyoming to prioritize community safety.”, ”Awareness and preparation among employers for the impending COVID-19 vaccine mandate.”, ”Against federal COVID-19 vaccine and mask mandates, framing them as overreach and questioning their constitutionality.”, ”Resistance against federal vaccine mandates, emphasizing personal choice and legal action to uphold individual rights.”, ”Vaccine documentation management, vaccine refusal, job consequences for non-compliance, and HIPAA implications in vaccine proof requests.”, ”Advocates for personal choice and against federal vaccine mandates, viewing them as infringements on civil liberties and rights.”
@@
VaccineReligion ”Using religion as a means to exempt individuals from mandatory COVID-19 vaccination requirements.”, ”Incorporating religious teachings into actions promoting community well-being, including vaccination and pandemic safety measures.”, ”Vaccination as an act of love and protection towards neighbors, aligning it with Christian values.”, ”Seeking religious guidance and perspective on COVID, vaccination, and related moral concerns with Pastor Bunjee Garrett.”, ”Religious communities to participate in public health by supporting vaccination as an act of neighborly love and care.”
@@
VaccineEfficacy ”The safety and efficacy of the Moderna COVID-19 vaccine, but acknowledges instances of post-vaccination infections.”, ”Getting vaccinated, highlighting the vaccine’s effectiveness in significantly reducing infection, severity, and mortality from the Delta variant of COVID-19.”, ”Strongly advocates for COVID-19 vaccination, highlighting its safety, efficacy, and crucial role in preventing severe illness and ending the pandemic.”, ”The prioritization strategy for COVID-19 vaccinations in certain highly infected North Carolina counties, addressing vaccine efficacy concerns.”, ”Promoting vaccination as an effective measure to protect individuals and families from COVID-19, with community endorsement.”, ”Highlights that vaccinations, including boosters, are highly effective in preventing severe illness, hospitalization, and death from COVID-19.”, ”Advocates for vaccination and continued safety measures to protect against the transmissible and potentially severe Delta variant.”
@@
VaccineDevelopment ”Trust in COVID-19 vaccines, highlighting their development from decades-old, scientifically robust messenger RNA technology.”, ”Confidence in the Pfizer-BioNTech COVID-19 vaccine, emphasizing its emergency use authorization and pending full FDA approval.”, ”Understanding the distinct characteristics and allocation strategies of the newly approved Johnson & Johnson COVID-19 vaccine.”
@@
CovidPlan ”Staying informed about regional COVID-19 vaccine distribution plans and understanding challenges in meeting demand.”, ”Recognizing and supporting the American Rescue Plan and lawmakers’ efforts in providing pandemic relief.”, ”Investigating and addressing the apparent inefficiencies in the COVID-19 vaccine distribution in Illinois.”, ”Commending the collaborative development of COVID-19 vaccines and treatments supported by the American Rescue Plan.”, ”Providing COVID-19 vaccinations to residents through the #VaxMobile program in different locations throughout the township.”
@@
VaccineMisinformation ”Correcting vaccine misinformation and encouraging vaccine participation through expert-led educational conversations.”, ”Debunking vaccine myths and empowering people with credible and accurate information about COVID-19 vaccines.”, ”Holding news outlets accountable for spreading potentially harmful misinformation about Covid-19 vaccines.”, ”Clear communication and spreading accurate information about COVID-19 vaccines to counter misinformation.”, ”Utilizing virtual town hall meetings to combat vaccine misinformation by providing accurate, expert information.
@@
NaturalImmunity ”Changes in herd immunity goals due to the Delta variant.”, ”Natural immunity against COVID-19 could be better than vaccine-induced immunity”, ”Boosting the immune system naturally with supplements as an alternative approach to handling COVID-19.”
A.4 Human Validation
Three NLP and CSS researchers, age range $30$-$45$, $2$ males and $1$ female, did two $1$-hour sessions to validate whether the generated and refined talking points as well as to annotate whether $ads\rightarrow arguments$ mapping is correct or not. The annotators included advanced graduate students and faculty. Our study was approved by the Institutional Review Board (IRB), and we followed their protocols.

### v3

@@ A.8 Demographic Analysis
CustBasedAltEnergy "Customer engagement in alternative energy solutions, emphasizing renewable and nuclear options to combat climate change and promote sustainability.", "Public support and awareness of nuclear energy as a sustainable, carbon-free alternative energy solution in the U.S.", "Community engagement and support for wind energy, highlighting its benefits for jobs, rural economies, and sustainable development.", "Community engagement in renewable energy through a cost-effective, equipment-free Community Solar program with enrollment incentives.", "Homeowner participation in solar energy trials to foster renewable energy adoption and reduce individual energy costs.", "Embracing solar energy for its broad benefits, focusing on economic growth, environmental protection, and enhanced accessibility.", "The adoption and support of utility-scale solar projects to supply clean energy to the grid for consumers.", "Embracing wind and nuclear power as key to Illinois’ energy strategy, highlighting job creation and environmental benefits.", "Adopting biofuels and innovative fuel technologies to reduce emissions and promote sustainability in transportation for consumers.", "Community engagement and legal actions to promote sustainable development and energy practices at the local level.", "Adopting geothermal heating and cooling as a cost-effective, efficient, and environmentally friendly alternative energy solution for homeowners."
@@
FoodSecurity "Addressing climate change’s impact on agriculture, social equity, and women’s rights to ensure food security for all.", "Cross-sector collaboration, addressing food loss, engaging communities, and prioritizing food security initiatives to combat global hunger.", "Participation in the Ration Challenge to support emergency food aid and attending a forum to explore food security issues.", "Engagement with the Fruitful Communities initiative to combat climate change, promote sustainable agriculture, and enhance food security.", "Urgent humanitarian aid and community support to address immediate and long-term food security challenges in crisis-affected areas.", "Sustainable dietary changes, particularly reducing meat consumption, to improve environmental health and ensure long-term food security.", "Community support and mobilization to provide essential resources and aid, ensuring food security and health during crises.", "Combining sustainable agriculture, traditional knowledge, and community efforts to improve food security and environmental sustainability."
@@
EnrgyAffrd&SustainLeg "Legislative action in Virginia to ensure energy reliability and affordability, learning from Texas’ deregulation disaster.", "Legislative action to significantly lower energy costs, promoting cleaner and more reliable energy solutions in North Carolina.", "Legislative and policy efforts to enhance energy efficiency, grid reliability, and sustainability across various regions.", "Legislative supporting utility bill assistance, pollution reduction, energy-efficient building codes, and transition to pollution-free buildings for sustainability.", "Policy-driven support for renewable energy, affordability, emissions reduction, and sustainable development, emphasizing community involvement and legislative backing.", "Legislative action to ensure utility service protection, improve energy infrastructure resilience, and address affordability amid extreme weather and economic challenges.", "Transitioning to sustainable energy sources, enhancing security for clean energy businesses, and supporting policies for cleaner, affordable energy solutions.", "Leveraging legislation and community programs to provide financial assistance for utility bills, ensuring energy affordability for all residents."
@@
EcofrndlyConsmChoice "Making eco-friendly consumer choices by purchasing sustainable apparel and actively contributing to environmental conservation efforts.", "Choosing eco-friendly, sustainable laundry solutions like Earthbreeze to reduce environmental impact and support ethical consumption.", "Adopting home water treatment systems to reduce reliance on single-use plastic bottles, promoting environmental sustainability and health.", "Making eco-friendly consumer choices by using Cleanyst’s plant-based products to reduce plastic waste and carbon emissions.", "Empowering communities and individuals through education and involvement in sustainable practices to combat climate change effectively.", "Embracing sustainable living through eco-friendly consumer choices, corporate sustainability efforts, and supporting green energy and conservation initiatives.", "Choosing sustainable, skin-friendly products like bamboo bandages as a simple yet effective way to support eco-friendly consumer choices.", "Supporting local leadership and community initiatives that promote sustainability, equitable growth, and eco-friendly consumer practices in urban development."
@@
PlasticWste&EnvImpact "Recognizing plastic’s sustainability benefits, urging scientific studies and informed policies to optimize its use and minimize environmental impact.", "Reducing plastic waste through reusable products, supporting cleanup efforts, and embracing innovative solutions to protect marine ecosystems.", "Using pyrolysis technology to recycle waste plastic, supporting a circular economy and environmental protection efforts.", "Urgent collective action, habit changes, and support for organizations fighting plastic pollution to protect oceans and marine life.", "Eengaging brands towards Plastic Neutrality and supports innovative recycling technologies to combat the global plastic waste crisis.", "Daily action against plastic waste, supporting legislation, and making conscious choices in food consumption and waste management.", "Responsible waste management, recycling efforts, and addressing illegal activities to mitigate environmental contamination and promote sustainability."
@@
PromtSustainTrnsport "Embracing sustainable transportation and clean energy through EV test drives and educational events on innovative energy solutions.", "Attending Drive Electric Earth Day events to learn about and embrace the benefits of electric vehicles for a sustainable future.", "Adopting cycling and supporting policies for electric school buses and a bold climate bill for sustainable transportation.", "Urgent investment in electric vehicles and legislative support to reduce emissions and create union jobs.", "Adopting electric vehicles for their environmental benefits, cost savings, and the expanding availability of models.", "Adopting clean car standards and investing in EV infrastructure to improve air quality, health, and combat climate change.", "Significant investment in public transportation electrification and infrastructure to support climate goals, community benefits, and union job creation.", "Adopting and promoting sustainable transportation solutions and technologies to improve air quality and address environmental challenges.", "Infrastructure investments and community initiatives that reduce emissions and promote sustainable transportation as essential to combating climate change.", "Adopting biking, walking, electric vehicles, and public transit as key strategies to reduce emissions and improve air quality.", "Shifting to public and active transportation, electric vehicles, and renewable energy to reduce emissions and enhance sustainability."
@@
WaterMngmnt&Sustain "Stricter regulations on coal ash disposal to protect water quality and public health in Georgia.", "Supporting legislative efforts to protect and restore Colorado’s rivers for economic growth, climate resilience, and agricultural preservation.", "Exploring and adopting successful global strategies to address water scarcity and ensure sustainable water management in rapidly growing states.", "Sustainable water management through refurbishing existing facilities and adopting eco-friendly water solutions to protect health and the environment.", "Proactive water management strategies, infrastructure investment, and community involvement to ensure sustainable water access and conservation efforts.", "Responsible disposal of hazardous wastes like oils and grease to protect water sources and promote environmental sustainability.", "Innovative financing, policy reforms, and sustainable practices to ensure water security and resilience against climate change impacts.", "Active engagement and informed leadership in water board governance to ensure sustainable water management and protection."
@@
GovDistrust "Accountability and reform in the Biden administration, criticizing its policies for undermining national stability and public trust.", "Against government control of the prescription drug market, warning of hindered medical innovation and reduced access to cures.", "Skepticism towards government actions and legislation, emphasizing the need for transparency and resistance to perceived overreach.", "Skepticism towards the Biden administration’s handling of COVID-19, immigration, and economic policies.", "Questions the efficacy and motives of government policies on immigration, economy, and public health measures."
@@
GovTrust "Recognizing and appreciating government efforts and collaboration in pandemic response, emphasizing trust in legislative support for public health.", "Recognizing the Biden administration’s efforts to address key national and global issues, building trust through proactive policies and actions.", "Political engagement and trust through transparency, community involvement, and proactive governance initiatives.", "Building trust through community engagement, open dialogue, and collaboration between citizens and their representatives.", "President Biden’s commitment to international accords, public health, economic recovery, and progressive policy achievements to build government trust."
@@
VaccineRollout "The widespread and accessible COVID-19 vaccination and testing services, emphasizing no-appointment-necessary clinics to increase public participation.", "Utilizing digital platforms to access up-to-date information on COVID-19 vaccine availability and registration across different locations.", "Easy and accessible COVID-19 vaccination for all eligible individuals, emphasizing convenience and inclusivity in public health efforts.", "The organized distribution of COVID-19 vaccines across Texas, highlighting state efforts to enhance accessibility and efficiency in vaccination.", "The importance of local news in keeping communities informed about the evolving COVID-19 vaccine rollout and registration processes.", "Accessible vaccine clinics, incentive programs, and safety protocols to enhance vaccine rollout at the local level."
@@
VaccineSymptom "Emphasizes the safety and effectiveness of COVID-19 vaccines despite mild side effects, advocating for vaccination to protect against the virus.", "Vigilance and thorough investigation into the rare but serious side effects like blood clots associated with the J&J vaccine.", "Careful monitoring and transparent communication about potential vaccine side effects to ensure public safety and confidence.", "Informed vaccine choices, awareness of rare side effects, and the importance of continued vaccination for public health safety.", "Awareness and transparency regarding the range and prevalence of COVID-19 vaccine-related symptoms and effects.", "Advocates for booster shots to enhance immunity, addressing vaccine safety and mild side effects experienced post-vaccination.", "Advocates for the safety of COVID-19 vaccines for children, emphasizing mild side effects and community protection through vaccination."
@@
VaccineEquity "Against price control legislation, emphasizing its potential to hinder vaccine development and equitable access to life-saving COVID-19 vaccinations.", "Global vaccine equity, emphasizing free and universal access to COVID-19 vaccines to end the pandemic for all.", "Overcoming barriers to ensure equitable COVID-19 vaccine access for underserved communities and prioritizing vulnerable populations.", "Investigating and improving the distribution and administration process to ensure equitable vaccine access in Illinois.", "Equitable access to COVID-19 vaccines, addressing eligibility, distribution fairness.", "Donating funds that will be matched to multiply impact, promoting fair global access to vaccines.", "Efforts and challenges in equitable vaccine distribution and access for seniors across various counties."
@@
VaccineStatus "Raising awareness about the low vaccination rates in Central Texas counties, as indicated by state data.", "Staying informed about COVID-19 vaccination updates and initiatives through trusted local news apps in respective counties.", "Accessing COVID-19 vaccine information and locations for vaccination through the SmartNews app for free.", "Highlights the concern over the high number of unvaccinated individuals in Central Texas counties according to state data."
@@
EncourageVaccination "Promoting vaccination by providing information, guidance, and incentives to encourage people to get vaccinated against COVID-19.", "Importance of COVID-19 vaccination by emphasizing its safety, efficacy, availability, and role in ending the pandemic.", "Engaging in community discussions and raising awareness about vaccination, especially within the LatinX community.", "Encourages COVID-19 vaccination and boosters for all ages to protect against the virus and maintain community health.", "Promoting COVID-19 vaccination for children and the community, emphasizing safety, accessibility, and the benefits of vaccination."
@@
VaccineMandate "Against COVID vaccine mandates and passports, urging support for personal choice and constitutional rights regarding vaccination.", "Awareness and updates on issues related to COVID-19 vaccine mandates and refusals in various states.", "COVID-19 vaccine mandates, including recent Supreme Court rulings and public opinion on federal mandates.", "Against Biden’s vaccine mandate, portraying it as government tyranny and urging resistance and financial support.", "Advocating for Covid-19 vaccine mandates for customers by Snowy Range Taxi in Wyoming to prioritize community safety.", "Awareness and preparation among employers for the impending COVID-19 vaccine mandate.", "Against federal COVID-19 vaccine and mask mandates, framing them as overreach and questioning their constitutionality.", "Resistance against federal vaccine mandates, emphasizing personal choice and legal action to uphold individual rights.", "Vaccine documentation management, vaccine refusal, job consequences for non-compliance, and HIPAA implications in vaccine proof requests.", "Advocates for personal choice and against federal vaccine mandates, viewing them as infringements on civil liberties and rights."
@@
VaccineReligion "Using religion as a means to exempt individuals from mandatory COVID-19 vaccination requirements.", "Incorporating religious teachings into actions promoting community well-being, including vaccination and pandemic safety measures.", "Vaccination as an act of love and protection towards neighbors, aligning it with Christian values.", "Seeking religious guidance and perspective on COVID, vaccination, and related moral concerns with Pastor Bunjee Garrett.", "Religious communities to participate in public health by supporting vaccination as an act of neighborly love and care."
@@
VaccineEfficacy "The safety and efficacy of the Moderna COVID-19 vaccine, but acknowledges instances of post-vaccination infections.", "Getting vaccinated, highlighting the vaccine’s effectiveness in significantly reducing infection, severity, and mortality from the Delta variant of COVID-19.", "Strongly advocates for COVID-19 vaccination, highlighting its safety, efficacy, and crucial role in preventing severe illness and ending the pandemic.", "The prioritization strategy for COVID-19 vaccinations in certain highly infected North Carolina counties, addressing vaccine efficacy concerns.", "Promoting vaccination as an effective measure to protect individuals and families from COVID-19, with community endorsement.", "Highlights that vaccinations, including boosters, are highly effective in preventing severe illness, hospitalization, and death from COVID-19.", "Advocates for vaccination and continued safety measures to protect against the transmissible and potentially severe Delta variant."
@@
VaccineDevelopment "Trust in COVID-19 vaccines, highlighting their development from decades-old, scientifically robust messenger RNA technology.", "Confidence in the Pfizer-BioNTech COVID-19 vaccine, emphasizing its emergency use authorization and pending full FDA approval.", "Understanding the distinct characteristics and allocation strategies of the newly approved Johnson & Johnson COVID-19 vaccine."
@@
CovidPlan "Staying informed about regional COVID-19 vaccine distribution plans and understanding challenges in meeting demand.", "Recognizing and supporting the American Rescue Plan and lawmakers’ efforts in providing pandemic relief.", "Investigating and addressing the apparent inefficiencies in the COVID-19 vaccine distribution in Illinois.", "Commending the collaborative development of COVID-19 vaccines and treatments supported by the American Rescue Plan.", "Providing COVID-19 vaccinations to residents through the #VaxMobile program in different locations throughout the township."
@@
VaccineMisinformation "Correcting vaccine misinformation and encouraging vaccine participation through expert-led educational conversations.", "Debunking vaccine myths and empowering people with credible and accurate information about COVID-19 vaccines.", "Holding news outlets accountable for spreading potentially harmful misinformation about Covid-19 vaccines.", "Clear communication and spreading accurate information about COVID-19 vaccines to counter misinformation.", "Utilizing virtual town hall meetings to combat vaccine misinformation by providing accurate, expert information.
@@
NaturalImmunity "Changes in herd immunity goals due to the Delta variant.", "Natural immunity against COVID-19 could be better than vaccine-induced immunity", "Boosting the immune system naturally with supplements as an alternative approach to handling COVID-19."